A Practical Version of the Generalized Lagrange Algorithm

Johannes Buchmann Max Jüntgen Michael Pohst

Abstract

We describe an implementation of the generalized Lagrange algorithm for computing units in algebraic number fields Nskip, together with extensive experimental data of the algorithm’s application (to all totally real quartic fields of discriminant below 60000). We also present an improved algorithm, with related experimental data.